A call option on a stock has a delta of 0.40 and a gamma of 0.10. A $0.10 rise in the price of the underlying stock will:

(a) reduce the price of the call option by $0.04

(b) reduce the delta of the call option by 0.04

(c) raise the price of an equivalent put option by $0.06

(d) raise the delta of the call option by 0.01.
Answer #1

The answer is

D. Raise the delta of call option by 0.01

Delta is change in option price with change in price of underlying security

The price will increase by 0.4*0.1 =$0.04

Delta will change by gamma*change in price

= 0.1*0.1 =0.01